Na2CO3 (s)  +  2HCl(g)   →  NaCl(s)  +  CO2 (g)  +  H2O(l)


      ∆H is -144.1 KJ. What is ∆E.


1
Expert's answer
2021-01-06T05:57:59-0500

"Na_2CO_{3(s)} + 2HCl_{(g)} \to NaCl_{(s)} + CO_{2(g)} + H_2O_{(l)} \qquad ∆H = -144.1 kJ"


"\u2206E = \u2206H - RT\u2206n"


"\u2206n = n_p - n_r = 1 - 2 = -1"

Assuming the reaction takes place at STP,

"RT\u2206n = 8.314\u00d7273\u00d7-1 = -2269.7J"

from the chemical reaction above, "\u2206H = -144.1kJ = -144,100J"



"\u2206E = -144100J - -(2269.7)J = -141830.3J"

"\therefore" ∆E = -141.8kJ